OpenStudy (anonymous):

passes through (3,-2) and has a slope -4 help please

OpenStudy (anonymous):

y=-4x-2

OpenStudy (the_fizicx99):

y-y1=m(x-x1) Plug in your values

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1\\ &(3\quad ,&-2) \end{array} \\\quad \\ slope = m= -4 \\ \quad \\ y-y_1=m(x-x_1)\qquad \textit{plug in your values and solve for "y"}\)